# Careless Whisper: Speech-to-Text Hallucination Harms

### v1

@@ 6. Calls to Action
First, Whisper API users should be made aware of the fact that their transcriptions could include hallucinations rendering the output inaccurate, or even harmful. Second, OpenAI should ensure the inclusion of more people with lived experience in designing systems such as Whisper. Third, OpenAI should work to calibrate Whisper’s default settings to reduce randomness in transcriptions. Fourth, the speech-to-text community should identify the causes of Whisper’s underlying hallucinations, and examine whether there are disproportionate negative effects on certain subpopulations. This knowledge can be used to inform practitioners on next steps to improve the Whisper model, such as more advanced modeling or data collection of specific types of speech acoustics leading to hallucinations.
@@ 7. Ethics Statement
Ethical considerations statement: The ethical concerns of this work mainly pertain to the data used for the analysis. The data was obtained from TalkBank’s AphasiaBank and therefore not directly collected by the researchers on our team. We had to rely on ethics considerations taken into account by TalkBank contributors when collecting the data. TalkBank has a code of ethics [^10] https://talkbank.org/share/ethics.html that enforces IRB-adherence and, additionally, data sharing norms and confidentiality practices, amongst others. By applying to TalkBank access, we adhered to this code of ethics. Furthermore, we opted out of data storage for all APIs tested, so TalkBank data for research purposes could not be repurposed for business use.
@@ A.1. Demographic Matching
The full data were subset to participants with provided demographic information. Then, matching on participant demographics was performed using the MatchIt package in R, using Mahalanobis matching (Rubin, 1980) and a caliper of 0.20. The matched covariates include gender, age, years of education, and English being a first language. The love plot below (including interaction variables) shows improved covariate balance post-matching. On the matched subset of 6,046 segments, we find that the share of segments yielding hallucinations remains significantly (p-value = 0.044) higher for aphasia speakers relative to control speakers (1.82% versus 1.16%, respectively).
@@ A.3. Regression Analysis
Table 3. By fitting a logistic regression conditioned on speaker demographics and audio segment attributes on various regression specifications (including matched subsets of audio segments per Section A.1), we still find that having longer non-vocal periods in an audio recording, and speaking with more words in the ground truth, tend to result in a significantly higher likelihood of a hallucinated Whisper transcription. The aphasia indicator coefficient is less significant than the regression conducted on the full set of audio data. Note that the matched datasets do not include covariates for English as a first language and/or vision loss due to the matched subsets consisting entirely of participants whose first language is English and who do not have vision loss.
@@
Has Aphasia 0.290 (0.204) $*{}^{*}$ (0.204) ${}^{***}$ (0.336) $*{}^{*}$ (0.248) ${}^{***}$ (0.439) ${}^{**}$ (0.443)
Non-vocal Duration (seconds) ${}^{***}$ (0.014) ${}^{***}$ (0.020) ${}^{**}$ (0.018) ${}^{**}$ (0.030) 0.044 (0.030)
Share of Duration Being Non-Vocal ${}^{**}$ (0.438)
Number of Words ${}^{***}$ (0.010) ${}^{***}$ (0.011) ${}^{***}$ (0.014) ${}^{***}$ (0.020)
Average Word Speed ${}^{***}$ (0.180) $*{}^{*}$ (0.362)
Is Female --0.022 (0.168) --0.017 (0.168) --0.053 (0.219) --0.054 (0.219) --0.141 (0.375) --0.118 (0.376)
Age 0.046 (0.043) 0.044 (0.043) 0.117 (0.081) 0.109 (0.080) 0.159 (0.172) 0.150 (0.171)
@@
African American --0.125 (0.468) --0.132 (0.469) 0.497 (0.597) 0.553 (0.598) 0.132 (1.048) 0.127 (1.049)
Other Race --0.295 (0.519) --0.281 (0.518) 0.108 (0.603) 0.185 (0.604) 0.429 (1.063) 0.473 (1.064)
Years of Education $*{}^{*}$ (0.033) $*{}^{*}$ (0.033) --0.068 (0.042) ${}^{**}$ (0.042) --0.073 (0.069) --0.087 (0.070)
No Hearing Loss --0.263 (1.035) --0.227 (1.035) 11.842 (549.080) 12.264 (534.209) 12.429 (841.479) 12.779 (817.992)
No Vision Loss 0.316 (1.023) 0.395 (1.026) 0.265 (1.032) 0.176 (1.034)
English is First Language 0.507 (1.027) 0.500 (1.028)
Constant ${}^{***}$ (2.157) ${}^{***}$ (2.175) --19.982 (549.086) --19.533 (534.215) --22.940 (841.495) --22.544 (818.009)
@@
Log Likelihood --781.440 --783.942 --457.257 --454.816 --156.581 --155.124
Akaike Inf. Crit. 1,588.879 1,593.883 938.515 933.631 335.161 332.247
Note: $*{}^{*}$p$<$0.1; ${}^{**}$p$<$0.05; ${}^{***}$p$<$0.01

### v2

@@ 6. Calls to Action
First, Whisper API users should be made aware of the fact that their transcriptions could include hallucinations rendering the output inaccurate, or even harmful. Second, OpenAI should ensure the inclusion of more people with lived experience in underserved communities (such individuals with aphasia) in designing systems such as Whisper. Third, OpenAI should work to calibrate Whisper’s default settings to reduce randomness in transcriptions. Fourth, the speech-to-text community should identify the causes of Whisper’s underlying hallucinations, and examine whether there are disproportionate negative effects on certain subpopulations. This knowledge can be used to inform practitioners on next steps to improve the Whisper model, such as more advanced modeling, or data collection of specific types of speech acoustics leading to hallucinations.
@@ 7. Ethics Statement
Ethical considerations statement: The ethical concerns of this work mainly pertain to the data used for the analysis. The data were obtained from TalkBank’s AphasiaBank and therefore not directly collected by the researchers on our team. We had to rely on ethics considerations taken into account by TalkBank contributors when collecting the data. TalkBank has a code of ethics [^10] https://talkbank.org/share/ethics.html that enforces IRB-adherence and, additionally, data sharing norms and confidentiality practices, amongst others. By applying to TalkBank access, we adhered to this code of ethics. Furthermore, we opted out of data storage for all APIs tested, so TalkBank data for research purposes could not be repurposed for business use.
@@ A.1. Demographic Matching
The full data were subset to participants with provided demographic information. Then, matching on participant demographics was performed using the MatchIt package in R, using Mahalanobis matching (Rubin, 1980) and a caliper of 0.20. The matched covariates include gender, age, race, years of education, and English being a first language. The love plot below (including interaction variables) shows improved covariate balance post-matching. On the matched subset of 6,046 segments, we find that the share of segments yielding hallucinations remains significantly (p-value = 0.044) higher for aphasia speakers relative to control speakers (1.82% versus 1.16%, respectively).
@@ A.3. Regression Analysis
Table 3. By fitting a logistic regression conditioned on speaker demographics and audio segment attributes on various regression specifications (including matched subsets of audio segments per Section A.1), we continue to find that having longer non-vocal periods in an audio recording, being aphasic, and speaking with more words (or greater speed) in the ground truth, tend to result in a higher likelihood of a hallucinated Whisper transcription (with varying levels of significance). Matched datasets do not include covariates for English as a first language and/or vision loss due to the matched subsets consisting entirely of participants whose first language is English and who do not have vision loss.
@@
Has Aphasia 0.626∗∗ (0.257) 0.290 (0.204) 0.917∗∗∗ (0.336) 0.517∗∗ (0.251) 1.188∗∗∗ (0.439) 0.947∗∗ (0.444)
Non-vocal Duration (seconds) 0.064∗∗∗ (0.015) 0.050∗∗∗ (0.014) 0.053∗∗∗ (0.020) 0.065∗∗ (0.030)
Average Word Speed 0.463∗∗∗ (0.128) 0.532∗∗∗ (0.180) 0.661 (0.362)
Share of Duration Being Non-Vocal 0.798 (0.556) --0.158 (0.992)
Number of Words 0.051∗∗∗ (0.010) 0.058∗∗∗ (0.015) 0.053∗∗ (0.022)
Is Female --0.019 (0.168) --0.022 (0.168) --0.053 (0.219) --0.052 (0.219) --0.141 (0.375) --0.098 (0.375)
Age 0.045 (0.043) 0.046 (0.043) 0.117 (0.081) 0.104 (0.079) 0.159 (0.172) 0.142 (0.169)
@@
African American --0.174 (0.468) --0.125 (0.468) 0.497 (0.597) 0.564 (0.598) 0.132 (1.048) 0.136 (1.049)
Other Race --0.311 (0.518) --0.295 (0.519) 0.108 (0.603) 0.211 (0.601) 0.429 (1.063) 0.528 (1.072)
Years of Education --0.048 (0.033) --0.061 (0.033) --0.068 (0.042) --0.081 (0.042) --0.073 (0.069) --0.092 (0.070)
No Hearing Loss --0.368 (1.037) --0.263 (1.035) 11.842 (549.080) 12.335 (533.761) 12.429 (841.479) 12.806 (810.943)
No Vision Loss 0.418 (1.023) 0.316 (1.023) 0.265 (1.032) 0.249 (1.038)
English is First Language 0.442 (1.024) 0.507 (1.027)
Constant --6.492∗∗∗ (2.169) --6.191∗∗∗ (2.157) --19.982 (549.086) --19.683 (533.767) --22.940 (841.495) --21.995 (810.960)
@@
Log Likelihood --785.562 --781.440 --457.257 --455.642 --156.581 --155.835
Akaike Inf. Crit. 1,597.125 1,588.879 938.515 935.284 335.161 333.670
Note: p<<<0.1; ∗∗p<<<0.05; ∗∗∗p<<<0.01